the [] of pure quinine, an antimalarial drug, is -165. If a solution contains75% quinine and...

the [\alpha] of pure quinine, an antimalarial drug, is -165. If a solution contains75% quinine and 25% of its enantiomer (ee= 50%), what is [\alpha] for the solution?

[\alpha] = ?
Answer #1

Given,

[\alpha] of pure quinine = -165

=> [\alpha] of its enantiomer = 165

Fraction of quinine in solution = 0.75

Fraction of enantiomer in solution = 0.25

=> [\alpha] for the solution = 0.75 x -165 + 0.25 x 165 = -82.5